# Rudyak’s conjecture for lower dimensional 1-connected manifolds

### v1

@@ Abstract.
Rudyak’s conjecture states that for any degree one map $f:M\to N$ between oriented closed manifolds there is the inequality $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)\geq\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)$ for the Lusternik-Shnirelmann category. We prove the Rudyak’s conjecture for $n$-dimensional simply connected manifolds for $n\leq 8$.
@@ 1. Introduction
In [17] Rudyak proved his conjecture with some restrictions on dimension, connectivity and the LS-category of the domain. It turned out that his conjecture is quite difficult even in the case of the most simple maps of degree one $f:M\#N\to N$, the collapsing of a summand in the connected sum M#NM\#N. That special case of Rudyak’s conjecture was settled in [6] where we proved the formula
$\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M\#N)=\max\{\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M),\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)\}.$
Perhaps the most interesting and still a challenging case of Rudyak’s conjecture is the case of birational morphism $f:M^{\prime}\to M$ between nonsingular projective varieties. We note when $M^{\prime}$ is a blowup of $M$ at a point the Rudyak’s conjecture for $f$ follows from the above formula.
In [4] there was an attempt to construct a counterexample to the Rudyak conjecture by means of Iwase’s manifolds. Using Iwase’s examples the first author constructed a map of degree one $f:M\to N$ between 32-dimensional simply-connected manifolds which would be a counter-example to Rudyak’s conjecture if $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M_{2}\times M_{3})\geq 5$. We recall that $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M_{2})=\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M_{3})=3$. The classical inequality [2]
@@
In any case we believe in Rudyak’s conjecture for lower dimensional manifolds. In [19] Rudyak verified his conjecture for manifolds of dimension $\leq 4$. In this paper we prove Rudyak’s conjecture for simply connected manifolds of dimension $\leq 8$ (Theorem 5.6).
@@ 2.1. Properties of maps of degree one
2.2. Whitehead Theorem
For nice spaces, such as CW complexes, it is an easy observation that $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}X\leq\dim X$. D.P. Grossmann [8] and independently G.W. Whitehead [21][2] proved that for simply connected CW complexes $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}X\leq\dim X/2$. Moreover, for $k$-connected CW complexes
@@ 3. Reduction to the 2 vs 3 case
All manifolds in the paper are considered to be smooth.
@@
By the Whitehead theorem $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}N\leq 2$. If $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)=1$ then $M$ is a homotopy sphere. By Proposition 2.1 $N$ is a homology sphere. Since $N$ is simply connected, it is a homotopy sphere. Hence $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)=1$. So, always $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)\geq\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)$. ∎
@@
By Whitehead’s theorem the LS-category of simply connected manifolds of dimension $\leq 7$ does not exceed 3. So, we have to consider only the case of $n=8$ where $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M,\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}N\leq 4$. Assume that $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M=3$ and $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}N=4$. Then there is no section of the Ganea fibration $p_{3}:G_{3}(N)\to N$. The fiber $F$ of this fibration is homotopy equivalent to a 6-connected space $\ast^{4}\Omega(N)$. The primary obstruction $\kappa\in H^{8}(N;\pi_{7}(F))$ to a section of $p_{3}$ is the only obstruction. Since the primary obstruction is natural, $f^{*}(\kappa)$ is the primary obstruction to the lift of $f$ with respect to $p_{3}$. By Proposition 2.1 $f^{*}(\kappa)\neq 0$. On the other hand, the inequality $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M\leq 3$ implies that such a lift does exist. Hence the obstruction is zero, $f^{*}(\kappa)=0$, which brings a contradiction. ∎
@@
(a) Suppose that $M^{\prime}$ is obtained from a simply connected $n$-manifold $M$ with $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M\leq 2$ by a surgery in dimensions $k\leq 4$. Then $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M^{\prime}\leq 2$.
(b) Suppose that $M^{\prime}$ is obtained from a connected manifold $M$ with $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M\leq 2$ by a surgery in dimension 1. Then $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M^{\prime}\leq 2$.
@@
(a) The inequality $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M\leq 2$ implies that there is a section $\sigma:M\to G_{2}(M)$ of the Ganea fibration $p_{M}:G_{2}(M)\to M$. Let $X=M\cup D^{k+1}$, $k\leq 4$, be obtained from $M$ by attaching a $(k+1)$-ball along an embedded $k$-sphere $S=S^{k}$. We show that $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}X\leq 2$. Note that the fiber of the Ganea fibration $p_{X}:G_{2}(X)\to X$ is 4-connected. Moreover, for any contractible subset $C\subset X$, the preimage $p_{X}^{-1}(C)$ is 4-connected. Let $p_{X}^{M}:G^{M}_{2}(X)\to M$ be the restriction of $p_{X}$ over $M$. Clearly $p_{X}^{M}$ contains $p_{M}$ as a subfibration. Then the restriction of $\sigma$ to $S$ is a null-homotopic map $\sigma|_{S}:S^{k}\to p_{X}^{-1}(D^{k+1})$. Let $\sigma^{\prime}:D^{k+1}\to p_{X}^{-1}(D^{k+1})$ be an extension. Then the union $\sigma\cup\sigma^{\prime}$ is a homotopy section of $p_{X}$. Then the Homotopy Lifting Property produces an honest section of $p_{X}$.
We note that the manifold $M^{\prime}$ admits a map $f:M^{\prime}\to X$ with fibers $S^{n-k-1}$. Thus, being a $UV^{n-k-2}$-map between CW-complexes, $f$ is (nk1)(n-k-1)-equivalence. Then $\Omega(f):\Omega(M^{\prime})\to\Omega(X)$ is an (nk2)(n-k-2)-equivalence. By Proposition 2.3 the induced map
@@
between the fibers of $G_{2}(M^{\prime})$ and $G_{2}(X)$ is a (nk+4)(n-k+4)-equivalence. This implies that the map $q:G_{2}(M^{\prime})\to Z$ to the pull-back in
@@
is a (nk+4)(n-k+4)-equivalence. Now a section $\sigma:X\to G_{2}(X)$ defines a section $s:M^{\prime}\to Z$. For $k\leq 4$ the map $q$ is a $n$-equivalence, therefore the map $s$ can be lifted with respect to $q$. Thus, $p_{M^{\prime}}$ admits a section and, hence, $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M^{\prime}\leq 2$.
(b) In general case the fiber $\ast^{3}\Omega(X)$ of $p_{X}:G_{2}(X)\to X$ is 1-connected and the map $q$ is a (nk+1)(n-k+1)-equivalence. Thus the argument works for $k=1$. ∎
4. Non-spin case
Theorem 4.1.
Every non-spin $n$-manifold $M$, $n\geq 7$, has $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M\geq 3$.
Every non-orientable $n$-manifold $M$, $n\geq 3$, has $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M\geq 3$.
Proof.
Let $\nu:M\to BSO$ be a map classifying the stable normal bundle on $M$. Suppose that $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M\leq 2$. We note that the kernel
$K=\ker\{\pi_{2}(M)\to\pi_{2}(BSO)=\mathbb{Z}_{2}\}$
is finitely generated and every 2-sphere $S\subset M$ representing an element of $K$ has trivial normal bundle. We perform the 2-surgery on $M$ to kill $K$ and obtain a new manifold $M^{\prime}$ such that its classifying map $\nu^{\prime}:M^{\prime}\to BSO$ is a 3-equivalence. By Proposition 3.3 $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M^{\prime}\leq 2$.
Next we do the 3-surgery on $M^{\prime}$ to kill $\pi_{3}(M^{\prime})$. We can do it, since $\dim M^{\prime}\geq 7$ and $\pi_{3}(BSO)=0$. As the result we obtain a manifold $M^{\prime\prime}$ whose classifying map $\nu^{\prime\prime}:M^{\prime\prime}\to BSO$ is a 4-equivalence. By Proposition 3.3 $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M^{\prime\prime}\leq 2$. Since $\nu^{\prime\prime}$ is a 4-equivalence, there is a homotopy lift $s:BSO^{(4)}\to M^{\prime\prime}$, i.e., such a map $s$ that the composition $\nu^{\prime\prime}s$ is homotopic to the inclusion $j:BSO^{(4)}\to BSO$. Let $w_{2}\in H^{2}(BSO;\mathbb{Z}_{2})$ be the 2nd Stiefel-Whitney class. Note that $w_{2}^{2}\neq 0$ [16]. The homotopy commutative triangle implies for the mod 2 cohomology
$0\neq j^{*}(w_{2}^{2})=s^{*}(\nu^{\prime\prime})^{*}(w_{2}^{2}).$
Hence $w_{2}(M)^{2}=((\nu^{\prime\prime})^{*}(w_{2}))^{2}=(\nu^{\prime\prime})^{*}(w_{2}^{2})\neq 0$. By Poincare duality there is a cohomology class $b\in H^{n-4}(M^{\prime\prime};\mathbb{Z}_{2})$ such that $w_{2}(M)^{2}b\neq 0$. Thus, the cup-length of $M^{\prime\prime}$ is at least 3, hence, $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M^{\prime\prime}\geq 3$. Contradiction.
The non-orientable case can be done with the use of the 1st Stiefel-Whitney class by doing 1-surgery on $\nu:M\to BO$ to get a 2-equivalence by application of Proposition 3.3(b). Note that the kernel $\ker\{\nu_{*}:\pi_{1}(M)\to\pi_{1}(BO)=\mathbb{Z}_{2}\}$ is finitely generated as an index two subgroup of a finitely generated group. ∎
@@ 5. Main result
By Whitehead Theorem $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)\leq 3$. Let us assume that $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)=2$. If $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)=1$ then $M$ is a homotopy sphere, so $N$ is a homotopy sphere in view of $\deg f=1$, hence $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)=1$, which is a contradiction.
@@
Let $f:M\to N$ be a map of degree 1 between $n$-dimensional simply connected manifolds that induces an isomorphism of homotopy groups $f_{*}:\pi_{i}(M)\to\pi_{i}(N)$ for $i\leq n-5$. Then
$\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)\geq\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N).$
@@
For $n=6$ this is the statement that Rudyak’s conjecture holds true for simply connected 6-dimensional manifolds which is proven in Corollary 5.3.
For $n\geq 9$ the conditions of the Lemma imply that $f$ is a homotopy equivalence, then $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)=\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)$.
Thus, the Lemma is about dimension 7 and 8. In this case Proposition 3.3 implies the inequality $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)\geq\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)$ when $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)\geq 3$. Also we can exclude the case when $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)=1$.
So, we consider the case $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)=2$ .
Now, since $M$ and $N$ are simply connected, we have $\pi_{2}(M)$ isomorphic to $H_{2}(M)$ by the Hurewicz theorem. Similarly, $\pi_{2}(N)$ is isomorphic to $H_{2}(N)$. Hence, the map $f$ induces an isomorphism $f_{*}:H_{2}(M)\to H_{2}(N)$. By the Poincare duality, $f^{*}:H^{5}(N)\to H^{5}(M)$ is an isomorphism.
Since $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)=2$, there is a section of the Ganea fibration
$p_{2}^{M}:G_{2}(M)\to M.$
The naturality of the Ganea construction implies that there a lift $\phi:M\to G_{2}(N)$ of $f$ with respect to $p_{2}^{N}$.
@@
Since $M$ and $N$ are simply connected, the condition of the lemma imply that $f_{*}:H_{i}(M)\to H_{i}(N)=\pi_{i}(N)$ are isomorphisms for $i\leq n-5$ by the classic Whitehead theorem. By the Poincare duality, $f^{*}:H^{i}(N)\to H^{i}(M)$ are isomorphisms for $i\geq 5$.
Note that the fiber $F=\ast^{5}\Omega N$ of the fibration $p=p_{2}^{N}:G_{2}(N)\to N$ is is the join product of 5 copies of connected space $\Omega N$ and hence it is $4$-connected. Therefore in the Moore-Postnikov tower for $p:G_{2}(N)\to N$ we have $N=Z_{1}=Z_{2}=Z_{3}=Z_{4}=Z_{5}$:
@@
By the Homotopy Lifting property of $q_{6}$, we can homotop the map $p_{7}\phi$ to a map $\psi:M\to Z_{7}$ such that $q_{6}\psi=s_{5}f$. Thus the map
@@
has two components $(s_{5}f,\psi_{1})$. Since $f$ induces an isomorphism of 2-dimensional homology groups, by the Poincare Duality it induces an isomorphism of 6-dimensional cohomology groups. Therefore, there is a map
@@
is null-homotopic. Hence $\kappa_{7}\psi$ is null-homotopic and the map $\kappa_{7}s^{\prime}s_{5}f$ is null-homotopic. This means that $f^{*}$ takes the cohomology class $\kappa_{7}s^{\prime}s_{5}]$ to zero. Since by Proposition 2.1 the homomorphism $f^{*}$ is injective, we obtain that $\kappa_{7}s^{\prime}s_{5}]=0$. Thus, the map $\kappa_{7}s^{\prime}s_{5}$ is null-homotopic and, hence, there is a lift $s_{7}:N\to Z_{8}$ of $s^{\prime}s_{5}$ with respect to $q_{7}$.
@@
A degree one map $f:M\to N$ of a $n$-dimensional, $n\geq 7$ simply connected spin manifold is bordant by means of surgery in dimension 2 and 3 to a a degree one map $f^{\prime}:M^{\prime}\to N$ that induces isomorphisms $f^{\prime}:\pi_{i}(M^{\prime})\to\pi_{i}(N)$ for $i\leq 3$.
@@
Note that $f_{*}:H_{2}(M)\to H_{2}(N)$ is surjective with a finitely generated kernel $K=kerf_{*}$ which is by the Hurewicz Theorem isomorphic to $ker\{f_{*}:\pi_{2}(M)\to\pi_{2}(N)\}$. Let $S_{1},\dots,S_{k}$ be 2-spheres smoothly embedded in $M$ that generate $K$. For each sphere $S_{i}$ its normal bundle $\nu_{i}$ is classified by a map $\nu_{i}:S_{i}\to BSO(5)$. The map $\nu_{i}$ is null-homotopic, since otherwise the evaluation of the Stiefel-Whitney class $w_{2}\in H^{2}(BSO(5);\mathbb{Z}_{2})$ on the image $(\nu_{i})_{*}([S_{i}])\in H_{2}(BSO(5);\mathbb{Z}_{2})$ would be nonzero. The latter would contradict to the spin condition on $M$. Thus, we can perform a surgery to obtain a bordism $W$ of $M$ to $M^{\prime}$ together with a map $\bar{f}:W\to N$. The restriction $f^{\prime}:M^{\prime}\to N$ of $\bar{f}$ to $M^{\prime}$ has degree one and $f^{\prime}$ induces an isomorphism of 2-dimensional homotopy groups.
By the above we may assume that $f:M\to N$ induces an isomorphism of 2-dimensional homotopy groups. Since $\pi_{3}(BSO(4))=0$, we obtain that the normal bundle of any embedded 3-sphere in $M^{\prime}$ is trivial. Hence we can perform surgery on generators of the kernel $ker\{f_{*}:\pi_{3}(M^{\prime})\to\pi_{3}(N)\}$ to get a bordism $W$ between $M$ and $M^{\prime}$ together with a map $\hat{f}:W\to N$ that agrees with $f$ on $M$ such that the restriction $f^{\prime}:M^{\prime}\to N$ of $\hat{f}$ to $M^{\prime}$ has degree one and it induces isomorphisms of homotopy groups of $\dim\leq 3$. ∎

### v2

@@ Abstract.
Rudyak’s conjecture states that for any degree one map $f:M\to N$ between oriented closed manifolds there is the inequality $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)\geq\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)$ for the Lusternik-Shnirelmann category. We prove the Rudyak’s conjecture for $n$-dimensional simply connected spin manifolds for $n\leq 8$.
@@ 1. Introduction
In [18] Rudyak proved his conjecture with some restrictions on dimension, connectivity and the LS-category of the domain. It turned out that his conjecture is quite difficult even in the case of the most simple maps of degree one $f:M\#N\to N$, the collapsing of a summand in the connected sum [7]. Perhaps the most interesting and still a challenging case of Rudyak’s conjecture is the case of birational morphism $f:M^{\prime}\to M$ between nonsingular projective varieties. We note when $M^{\prime}$ is a blowup of $M$ at a point the Rudyak’s conjecture for $f$ follows from the above formula.
Rudyak’s conjecture like many other natural conjectures on the LS-category could be false. In [5] there was an attempt to construct a counterexample to the Rudyak conjecture by means of Iwase’s manifolds. Using Iwase’s examples the first author constructed a map of degree one $f:M\to N$ between 32-dimensional simply-connected manifolds which would be a counter-example to Rudyak’s conjecture if $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M_{2}\times M_{3})\geq 5$. We recall that $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M_{2})=\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M_{3})=3$. The classical inequality [2]
@@
A different idea for a counterexample was suggested in [4].
In any case we believe in Rudyak’s conjecture for lower dimensional manifolds. In [20] Rudyak verified his conjecture for manifolds of dimension $\leq 4$. In this paper we prove Rudyak’s conjecture for simply connected spin manifolds of dimension $\leq 8$ (Theorem 4.6). At the end of the paper we give an example of non-spin simply connected 7-dimensional manifold. The existence of such an example does not allow us to extend our approach to non-spin manifolds.
@@ 2.1. Properties of maps of degree one
2.2. G. W. Whitehead Theorem
For nice spaces, such as CW complexes, it is an easy observation that $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}X\leq\dim X$. D.P. Grossmann [9] and independently G.W. Whitehead [24][2] proved that for simply connected CW complexes $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}X\leq\dim X/2$. Moreover, the following inequality holds true [2].
Theorem 2.2.
For $k$-connected CW complexes
@@ 2.3. Ganea’s approach to the LS-category
We recall that the LS-category of a map $f:Y\to X$ is the least integer $k$ such that $Y$ can be covered by $k+1$ open sets $U_{0},\dots,U_{k}$ such that the restrictions $f|_{U_{i}}$ are null-homotopic for all $i$. The above theorem can be extended to maps [4]:
Theorem 2.4.
[…]
@@ 3. Reduction to the 2 vs 3 case
All manifolds in the paper are considered to be connected, closed, orientable, and smooth.
@@
By Theorem 2.2 $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}N\leq 2$. If $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)=1$ then $M$ is a homotopy sphere. By Proposition 2.1 NN is a homology sphere. Since $N$ is simply connected, by the Hurewicz theorem it is a homotopy sphere. Then by the J. H. C. Whitehead theorem NN is homotopy equivalent to $S^{n}$. Hence $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)=1$. So, always $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)\geq\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)$. ∎
@@
By Theorem 2.2 the LS-category of simply connected manifolds of dimension $\leq 7$ does not exceed 3. So, we have to consider only the case of $n=8$ where $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M,\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}N\leq 4$. Assume that $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}N=4$. Then there is no section of the Ganea fibration $p_{3}:G_{3}(N)\to N$. The fiber $F$ of this fibration is homotopy equivalent to a 6-connected space $\ast^{4}\Omega(N)$. The primary obstruction $\kappa\in H^{8}(N;\pi_{7}(F))$ to a section of $p_{3}$ is the only obstruction. Since the primary obstruction is natural, $f^{*}(\kappa)$ is the primary obstruction to the lift of $f$ with respect to $p_{3}$. By Proposition 2.1 $f^{*}(\kappa)\neq 0$. Hence by Proposition 2.4 $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}f>3$. The inequality $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}f\leq\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M$ completes the proof. ∎
We recall that if a $n$-manifold $M^{\prime}$ is obtained from $M$ by a surgery then there is a bordism beetween them, i.e. $(n+1)$-dimenisonal manifold $W$ with boundary $\partial W=M\sqcup M^{\prime}$. If the surgery was in dimension $k$, then $W$ is homotopy equivalent to $M$ with some $(k+1)$-cells attached. If $M^{\prime}$ is obtained by a $k$-surgery from $M$, then $M$ is obtained by (nk1)(n-k-1)-surgery from $M^{\prime}$.
@@
(a) Suppose that $M^{\prime}$ is obtained by a $k$-surgery, $k\leq 4$, from a simply connected $n$-manifold $M$ with $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M\leq 2$. Then $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}W\leq 2$ where $W$ is a bordism of the surgery.
(b) Suppose that $M^{\prime}$ is obtained from a simply connected manifold $M$ with $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M\leq 2$ by a surgery in dimension 2. Then $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M^{\prime}\leq 2$.
@@
(a) The inequality $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M\leq 2$ implies that there is a section $\sigma:M\to G_{2}(M)$ of the Ganea fibration $p_{M}:G_{2}(M)\to M$. The bordism $W$ is homotopy equivalent to the space
$X=M\cup(\cup_{\alpha}D^{k+1}_{\alpha})$
obtained from $M$ by attaching a $(k+1)$-balls along disjoint embedded $k$-spheres $S^{k}_{\alpha}$. We show that $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}X\leq 2$. Note that the fiber of the Ganea fibration $p_{X}:G_{2}(X)\to X$ is 4-connected. Moreover, for any contractible subset $C\subset X$, the preimage $p_{X}^{-1}(C)$ is 4-connected. Let $p_{X}^{M}:G^{M}_{2}(X)\to M$ be the restriction of $p_{X}$ over $M$. Clearly $p_{X}^{M}$ contains $p_{M}$ as a subfibration. Then the restriction of $\sigma$ to $S_{\alpha}$ is a null-homotopic map $\sigma|_{\alpha}:S^{k}_{\alpha}\to p_{X}^{-1}(D^{k+1})$ for all $\alpha$. Let
$\sigma^{\prime}_{\alpha}:D^{k+1}_{\alpha}\to p_{X}^{-1}(D^{k+1}_{\alpha})$
be an extension. Then the union $\sigma\cup(\cup_{\alpha}\sigma^{\prime}_{\alpha}$ is a homotopy section of $p_{X}$. Then the Homotopy Lifting Property produces an honest section of $p_{X}$.
(b) We note that $W$ is homotopy equivalent to the space
$X^{\prime}=M^{\prime}\cup(\cup_{\alpha}D^{n-2}_{\alpha}).$
Hence the inclusion $j:M^{\prime}\to X^{\prime}$ is (n3)(n-3)-equivalence. Then $\Omega(j):\Omega(M^{\prime})\to\Omega(X^{\prime})$ is an (n4)(n-4)-equivalence. By Proposition 2.5 the induced map
@@
between the fibers of $G_{2}(M^{\prime})$ and $G_{2}(X^{\prime})$ is a (2+2+n4)(2+2+n-4)-equivalence. This implies that the map $q:G_{2}(M^{\prime})\to Z$ to the pull-back in
@@
is a $n$-equivalence. A section $\sigma:X^{\prime}\to G_{2}(X^{\prime})$ defines a section $s:M^{\prime}\to Z$. Since the map $q$ is a $n$-equivalence, the map $s$ can be lifted with respect to $q$. Thus, $p_{M^{\prime}}$ admits a section and, hence, by Theorem 2.3 $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M^{\prime}\leq 2$. ∎
@@ 4. Main result
By Theorem 2.2 $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)\leq 3$. Let us assume that $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)=2$. If $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(M)=1$ then $M$ is a homotopy sphere, so $N$ is a homotopy sphere in view of $\deg f=1$, hence $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)=1$, which is a contradiction.
@@
Let $f:M\to N$ be a map of degree 1 between $n$-dimensional simply connected manifolds where $n=7$ or $8$ that induces an isomorphism of homotopy groups $f_{*}:\pi_{i}(M)\to\pi_{i}(N)$ for $i\leq n-5$. Then $\max\{\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(f),2\}\geq\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(N)$.
@@
[…]
Now, since $M$ and $N$ are simply connected, we have $\pi_{2}(M)$ isomorphic to $H_{2}(M)$ and $\pi_{2}(N)$ is isomorphic to $H_{2}(N)$ by the Hurewicz theorem. Hence, the map $f$ induces an isomorphism $f_{*}:H_{2}(M)\to H_{2}(N)$.
Since $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}(f)\leq 2$, By Theorem 2.4 there is a lift $\phi:M\to G_{2}(N)$ of $f$ with respect to the Ganea fibration $p_{2}^{N}$,
@@
Since $M$ and $N$ are simply connected, the condition of the lemma imply that $f_{*}:H_{i}(M)\to H_{i}(N)=\pi_{i}(N)$ are isomorphisms for $i\leq n-5$ by the classic Whitehead theorem. Then by the Poincare duality, $f^{*}:H^{i}(N)\to H^{i}(M)$ are isomorphisms for $i\geq 5$.
Note that the fiber $F=\ast^{3}\Omega N$ of the fibration $p=p_{2}^{N}:G_{2}(N)\to N$ is is the join product of 3 copies of connected space $\Omega N$ and hence it is $4$-connected. Therefore in the Moore-Postnikov tower for $p:G_{2}(N)\to N$ we have $N=Z_{1}=Z_{2}=Z_{3}=Z_{4}=Z_{5}$:
@@
Since $p_{6}\phi$ is homotopic to $s_{5}f$ and $q_{6}p_{7}\phi=p_{6}\phi$, by the Homotopy Lifting property of $q_{6}$, we can homotop the map $p_{7}\phi$ to a map $\psi:M\to Z_{7}$ such that $q_{6}\psi=s_{5}f$. Thus the map
@@
has two components $(s_{5}f,\psi_{1})$. Since $f$ induces an isomorphism of 6-dimensional cohomology groups, there is a map
@@
is null-homotopic. Hence $\kappa_{7}\psi$ is null-homotopic and the map $\kappa_{7}s^{\prime}s_{5}f$ is null-homotopic. This means that $f^{*}$ takes the cohomology class $\kappa_{7}s^{\prime}s_{5}]$ to zero. Since the homomorphism $f^{*}$ is injective, we obtain that $\kappa_{7}s^{\prime}s_{5}]=0$. Thus, the map $\kappa_{7}s^{\prime}s_{5}$ is null-homotopic and, hence, there is a lift $s_{7}:N\to Z_{8}$ of $s^{\prime}s_{5}$ with respect to $q_{7}$.
@@
A degree one map $f:M\to N$ of a $n$-dimensional, $n=7,8$ simply connected spin manifold with $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M\leq 2$ is bordant by means of surgery in dimension 2 and 3 to a a degree one map $f^{\prime}:M^{\prime}\to N$ with $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}f^{\prime}\leq 2$ that induces isomorphisms $f^{\prime}:\pi_{i}(M^{\prime})\to\pi_{i}(N)$ for $i\leq n-5$.
@@
For $n=7,8$ by Proposition 2.1 the induced homomorphism $f_{*}:H_{2}(M)\to H_{2}(N)$ is surjective with a finitely generated kernel $K=kerf_{*}$ which is by the Hurewicz Theorem isomorphic to $ker\{f_{*}:\pi_{2}(M)\to\pi_{2}(N)\}$. Let $S_{1},\dots,S_{k}$ be disjoint 2-spheres smoothly embedded in $M$ that generate $K$. For each sphere $S_{i}$ its normal bundle $\nu_{i}$ is classified by a map $\nu_{i}:S_{i}\to BSO(n-2)$. The map $\nu_{i}$ is null-homotopic, since otherwise the evaluation of the Stiefel-Whitney class $w_{2}\in H^{2}(BSO(n-2);\mathbb{Z}_{2})$ on the image $(\nu_{i})_{*}([S_{i}])\in H_{2}(BSO(n-2);\mathbb{Z}_{2})$ would be nonzero. The latter would contradict to the spin condition on $M$. Thus, we can perform a surgery to obtain a bordism $W$ of $M$ to $M^{\prime}$ together with a map $\bar{f}:W\to N$. The restriction $f^{\prime}:M^{\prime}\to N$ of $\bar{f}$ to $M^{\prime}$ has degree one and $f^{\prime}$ induces an isomorphism of 2-dimensional homotopy groups. For $n=7$ by Proposition 3.3 (b) $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}M^{\prime}\leq 2$ and, hence, $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}f^{\prime}\leq 2$.
When $n=8$ by the above we may assume that $f:M\to N$ induces an isomorphism of 2-dimensional homotopy groups. Since $\pi_{3}(BSO(n-3))=0$, we obtain that the normal bundle of any embedded 3-sphere in $M^{\prime}$ is trivial. Hence we can perform surgery on generators of the kernel $ker\{f_{*}:\pi_{3}(M^{\prime})\to\pi_{3}(N)\}$ to get a bordism $W$ between $M$ and $M^{\prime}$ together with a map $\hat{f}:W\to N$ that agrees with $f$ on $M$ such that the restriction $f^{\prime}:M^{\prime}\to N$ of $\hat{f}$ to $M^{\prime}$ has degree one and it induces isomorphisms of homotopy groups of $\dim\leq 3$. By Proposition 3.3 (a) we obtain that $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}W\leq 2$. Hence $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}\hat{f}\leq 2$. By Theorem 2.4 $\hat{f}$ admits a lift $\hat{\phi}:W\to G_{2}(N)$ of $\hat{f}$ with respect to $p_{2}^{N}$. The restriction of $\hat{\phi}$ to $M^{\prime}$ is a lift of $f^{\prime}:M^{\prime}\to N$ with respect to $p^{N}_{2}$. Then by Theorem 2.4 $\operatorname{{\mbox{\rm cat${}_{\rm LS}$}}}f^{\prime}\leq 2$. ∎